Ryan and Nathan’s ages add up to 68 years. If you subtracted their ages, you would get 4. How old is Ryan if he is the older sibling?
Answer by ikleyn(52788)

.

Solve it this way

    R + (R-4) = 68.

    2R = 68 + 4 = 72

    R  = 72/2 = 36.    ANSWER

OR this way

    R + N = 68

    R - N =  4

-------------------  Add the equations and get the same answer

   2R     = 68 + 4 = 72

    R              = 72/2 = 36  years old.